OpenStudy (anonymous):

let , be two vectors in a vector space V. Show that , are linearly dependent if and only if one of the vectors is a scalar multiple of the other.

OpenStudy (anonymous):

I understand that it is clear the two vectors are linearly dependent, I'm just having trouble proving this algebraically.

OpenStudy (anonymous):

Why do it algebraically if you can plot them in MATLAB :-D

OpenStudy (anonymous):

oh.... can't be visualized using MATLAB :(

OpenStudy (across):

You can prove that two vectors are linearly dependent by computing their Wronskian.

OpenStudy (across):

... or equivalent thereof.

OpenStudy (anonymous):

really? Had no idea the wronskian was used aside from variation of parameters.

OpenStudy (anonymous):

something about it being 0 and whatnot

OpenStudy (jamesj):

It's much easier than that, isn't it? Two vectors v and w are linearly dependent, by definition if there exist _non-zero_ constants a and b such that av + bw = 0. Now from that, you should be able to prove very easily your result.

OpenStudy (anonymous):

That's what I had originally thought. I was then able to say that (a+kb) = 0. This doesn't prove anything, but I'm thinking it might be a basis for the nullspace?

OpenStudy (anonymous):

From there I figured I could say that the existence of the nullspace proves linear depence.

OpenStudy (across):

Let\[\vec{v}=\begin{pmatrix} v_1\\ v_2 \end{pmatrix},\]\[\vec{u}=\begin{pmatrix} u_1\\ u_2 \end{pmatrix},\]\[\vec{u}=c\vec{v},\exists c\in\mathbb{R}\setminus\{0\}.\]Then\[\begin{vmatrix} v_1 & u_1\\ v_2 & u_2 \end{vmatrix}=\begin{vmatrix} v_1 & cv_1\\ v_2 & cv_2 \end{vmatrix}=...\]

OpenStudy (anonymous):

james is right. dont make this problem way harder than it has to be >.<

OpenStudy (jamesj):

If v and w are Lin Dep, there exists nonzero constants a and b such that av + bw = 0 => v = (-b/a)w. Thus one is scalar multiple of the other. Conversely, suppose wlog that v = kw for some constant k. Then v + (-k)w = 0. Hence v and w and Lin Dep.

OpenStudy (anonymous):

I understand the first proof provided. It makes perfect sense. However, the second one makes less sense. I don't understand how changing the constant to some other arbitrary constant proves linear dependence.

OpenStudy (jamesj):

Let v and w be two vectors in a vector space V. Then let A be the statement: "v and w are linearly dependent" and let B be the statement "v can be written as a scalar multiple of w or vice versa" We want to show that A => B and B => A The proof of A => B is the first thing: If v and w are Lin Dep, there exists nonzero constants a and b such that av + bw = 0 => v = (-b/a)w. Thus one is scalar multiple of the other. The proof of the converse, B => A is now this: Suppose wlog (without loss of generality) that v = kw for some constant k. Then v + (-k)w = 0. Hence v and w and Lin Dep. ====== We write v = kw for some constant k because that's our hypothesis; we're starting with the statement B; this mathematical statement is what we know.

OpenStudy (anonymous):

OH! Right. Hah, that's quite obvious now. Thanks!